OpenStudy (anonymous):

Solve for x: -3(x + 3) = -3(x + 1) - 5.

OpenStudy (solomonzelman):

expand both sides, and add like terms on each side

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{-3(x + 3) = -3(x + 1) - 5}\) \(\normalsize\color{blue}{-3x -9= -3x -3 - 5}\) Now, I am going to highlight the like terms. \(\normalsize\color{blue}{\color{darkgoldenrod}{-3x} \color{red}{-9}= \color{darkgoldenrod}{-3x} \color{red}{-3}\color{red}{-5}}\)

OpenStudy (solomonzelman):

go for it.